OpenStudy (anonymous):

Calc 3 question! Please help!

OpenStudy (anonymous):

Evaluate \[\int\limits_{C}^{}(x^2+y^2)dx + 2xydy\]where C is positively (anti-clockwise) oriented boundary of the region \[0\le x \le1\]and \[2x^2 \le y \le 2x\]

OpenStudy (dan815):